OpenStudy (rock_mit182):

help with this one please ! The bob of a pendulum swings through an arc 24 centimeters long on its first swing. If each successive swing is approximately five-sixths the length of the preceding swing, use an infinite geometric series to approximate the total distance the bob travels.

OpenStudy (nurali):

Infinite geometric series. Sum = a/(1-r) Sum = 24/(1-(5/6)) = 24/1-6 = 144 cm

OpenStudy (rock_mit182):

thanks a lot !

OpenStudy (nurali):

My Pleasure.

OpenStudy (rock_mit182):

hey but the symbol would be sigma in "sum"

OpenStudy (rock_mit182):

\[\sum_{0}^{a}\] ?